# A Hausdorff-measure boundary element method for acoustic scattering by fractal screens

### v1

@@ 2.5 Function spaces on screens
For open sets $\Omega\subset\mathbb{R}^{n+1}$ (e.g. the exterior domain $D:=\mathbb{R}^{n+1}\setminus{\Gamma}$) we work with the classical Sobolev spaces [^9] Our notation follows [32]. Given an open set $\Omega\subset\mathbb{R}^{n+1}$, $W^{1}(\Omega)$ is the Sobolev space whose norm is defined “intrinsically”, via integrals over $\Omega$, while $H^{1}(\Omega)$ is defined (see §2.4) “extrinsically” as the set of restrictions to $\Omega$ of functions in $H^{1}(\mathbb{R}^{n+1})$. These spaces coincide if $\Omega$ is Lipschitz (e.g., [32, Theorem 3.16]), in particular if $\Omega=\mathbb{R}^{n+1}$, but not in general, in particular not, in general, for $\Omega=D$. $W^{1}(\Omega)$ and $W^{1}(\Omega,\Delta)$, normed by $\|u\|_{W^{1}(\Omega)}^{2}=\|u\|_{L_{2}(\Omega)}^{2}+\|\nabla u\|_{L_{2}(\Omega)}^{2}$ and $\|u\|_{W^{1}(\Omega,\Delta)}^{2}=\|u\|_{L_{2}(\Omega)}^{2}+\|\nabla u\|_{L_{2}(\Omega)}^{2}+\|\Delta u\|_{L_{2}(\Omega)}^{2}$ respectively, and their “local” versions $W^{1,{\rm loc}}(\Omega)$ and $W^{1,{\rm loc}}(\Omega,\Delta)$, defined as the sets of measurable functions on $\Omega$ whose restrictions to any bounded open $\Omega^{\prime}\subset\Omega$ are in $W^{1}(\Omega^{\prime})$ or $W^{1}(\Omega^{\prime},\Delta)$ respectively. We denote by $\gamma^{\pm}:W^{1}(U^{\pm})\to H^{1/2}(\Gamma_{\infty})$ and $\partial_{\mathrm{n}}^{\pm}:W^{1}(U^{\pm},\Delta)\to H^{-1/2}(\Gamma_{\infty})$ the standard Dirichlet and Neumann trace operators from the upper and lower half spaces $U^{\pm}:=\{x\in\mathbb{R}^{n+1},\pm x_{n+1}>0\}$ onto the hyperplane $\Gamma_{\infty}$, where the normal vector is assumed to point into $U^{+}$ in the case of the Neumann trace. Explicitly, the traces are the extension by density of $\gamma^{\pm}(u)(x):=\lim_{\begin{subarray}{c}x^{\prime}\to x\\ x^{\prime}\in U^{\pm}\end{subarray}}u(x^{\prime})$ and $\partial_{\mathrm{n}}^{\pm}u(x):=\lim_{\begin{subarray}{c}x^{\prime}\to x\\ x^{\prime}\in U^{\pm}\end{subarray}}\frac{\partial u}{\partial x_{n+1}}(x^{\prime})$ for $u\in C^{\infty}_{0}(\mathbb{R}^{n+1})|_{U^{\pm}}$ and $x\in\Gamma_{\infty}$. We note that if $u\in W^{1}(\mathbb{R}^{n+1})$ then $\gamma^{+}(u|_{U^{+}})=\gamma^{-}(u|_{U^{-}})$ [12, §2.1]. Finally, let $C^{\infty}_{0,\Gamma}$ denote the set of functions in $C^{\infty}_{0}(\mathbb{R}^{n+1})$ that equal one in a neighbourhood of $\Gamma$.
@@ 3 Wavelet decompositions
The spaces $\mathbb{H}^{t}(\Gamma)$ were defined in §2.4 for $\Gamma\subset\mathbb{R}^{n}$ a general $d$-set with $0<d\leq n$. In the case where $\Gamma$ is a disjoint IFS attractor (in the sense of §2.3, in which case, by Lemma 2.6, $d<n$), it was shown in [30] that, when $n-1<d={\rm dim_{H}}(\Gamma)<n$, the spaces $\mathbb{H}^{t}(\Gamma)$ can be characterised in terms of wavelet decompositions for t>0t>0.101010More precisely, [30] showed that the Besov spaces $B^{p,q}_{\alpha}(\Gamma)$, as defined in [30, §6] (and see Definition A.1 below), can be characterised in this way for $\alpha>0$ and $1\leq p,q\leq\infty$, provided $\Gamma$ preserves Markov’s inequality, which holds in particular (see Remark A.6) if d>n1d>n-1. We show in Appendix A (see Corollary A.4 and Remark A.6) that, for $t>0$, $\mathbb{H}^{t}(\Gamma)=B^{2,2}_{t}(\Gamma)$ with equivalence of norms if $\Gamma$ is a $d$-set with d>n1d>n-1, so that this characterisation carries over to our trace spaces $\mathbb{H}^{t}(\Gamma)$. This, more precisely an extension of this characterisation to negative $t$, will be central to our BEM convergence analysis later. In the current section we recap the notation and main results from [30] that we will need, initially assuming only that $n-1<d\leq n$ and that the OSC is satisfied.
@@
These bases are hierarchical, in the sense that the basis for $\mathbb{W}_{\ell+1}$ contains that for $\mathbb{W}_{\ell}$. Furthermore, as noted in [30, p. 334] (and demonstrated in the proof of Theorem 5.1 below), elements of $\mathbb{L}_{2}(\Gamma)$ can be approximated arbitrarily well by elements of $\mathbb{W}_{\ell}$ as $\ell\to\infty$, which implies that
@@
The next result, which combines Theorems 1 and 2 in [30], provides a characterization of the space $\mathbb{H}^{t}(\Gamma)\subset\mathbb{L}_{2}(\Gamma)$ (introduced after (10)) for $n-1<d={\rm dim_{H}}(\Gamma)<n$ and $0<t<1$ in terms of the wavelet basis introduced above, under the assumption that $\Gamma$ is a disjoint IFS attractor. This assumption ensures that the piecewise-constant spaces $\mathbb{W}_{\ell}$ are contained in $\mathbb{H}^{t}(\Gamma)$ for all $t>0$ (see [30, Thm. 2]). In the statement of Theorem 3.1 the set $J_{\nu}$ is defined for $\nu\in\mathbb{Z}$ by
@@
Theorem 3.1 has the following important corollary, which is obtained by duality. In this corollary and subsequently (see, e.g., [13, Remark 3.8]), given an interval $\mathcal{I}\subset\mathbb{R}$ we will say that a collection of Hilbert spaces $\{H_{s}:s\in\mathcal{I}\}$, indexed by $\mathcal{I}$, is an interpolation scale if, for all $s,t\in\mathcal{I}$ and $0<\eta<1$, $(H_{s},H_{t})$ is a compatible couple (in the standard sense, e.g. [3, §2.3]) and if the interpolation space $(H_{s},H_{t})_{\eta}$ [^12] Here, and subsequently, $(H_{s},H_{t})_{\eta}$ denotes the standard complex interpolation space, $(H_{s},H_{t})_{[\eta]}$ in the notation of [3]; equivalently, the $K$- or $J$-method real interpolation spaces denoted $(H_{s},H_{t})_{\eta,2}$ in [3, 13], which are the same Hilbert spaces $(H_{s},H_{t})_{[\eta]}$, with equal norms, if the $K$- and $J$-methods are appropriately normalised (see [13, Remark 3.6] and [15]). coincides with $H_{\theta}$, for $\theta=(1-\eta)s+\eta t$, with equivalent norms. We will say that $\{H_{s}:s\in\mathcal{I}\}$ is an exact interpolation scale if, moreover, the norms of $(H_{s},H_{t})_{\eta}$ and $H_{\theta}$ coincide, for all $s,t\in\mathcal{I}$ and $0<\eta<1$.
@@
For (v) we note that by [13, Thm. 3.1] $\{\mathfrak{h}^{\tau}\}_{\tau\in\mathbb{R}}$ is an exact interpolation scale. The corresponding statement about $\{\mathbb{H}^{\tau}(\Gamma)\}_{-1<\tau<1}$, equipped with the norm $\|\cdot\|_{\tau}$, follows from Theorem 3.1 and parts (i)-(iii), combined with [^13] There is an inaccuracy in the statement of [13, Cor. 3.2]; the map $\mathcal{A}:\Sigma(\overline{H})\to\mathcal{Y}$ in that corollary needs to be injective as well as linear for the corollary to hold (see [15]). This injectivity follows automatically from the other conditions on $\mathcal{A}$ when (as in the application we make here) $H_{1}\subset H_{0}$, with continuous embedding, since then $\Sigma(\overline{H})=H_{0}$. [13, Cor. 3.2] and the fact that $\mathbb{H}^{0}(\Gamma)=\mathbb{L}_{2}(\Gamma)$ is unitarily isomorphic to $\mathfrak{h}^{0}$. Explicitly, in [13, Cor. 3.2], given $1<\tau_{0}<\tau_{1}<1$ we take $H_{j}=\mathbb{H}^{\tau_{j}}(\Gamma)$, $j=0,1$, $\mathcal{X}=\{0\}\cup\{(\nu,{\boldsymbol{m}},m):{\boldsymbol{m}}\in I_{\ell},\,\ell\in\mathbb{N}_{0},\,m\in\{1,\ldots,M-1\}\}$, $\mu$ to be the counting measure on $\mathcal{X}$, $\mathcal{A}$ to be the map taking $f\in\mathbb{H}^{\tau_{j}}(\Gamma)$ to the sequence $\boldsymbol{\beta}$ defined by (22) or (26) (as appropriate), and $w_{j}(0)=1$, $w_{j}((\nu,{\boldsymbol{m}},m))=2^{2\nu\tau_{j}}$. ∎
A basic interpolation result is that if $X_{0}\supset X_{1}$ and $Y_{0}\supset Y_{1}$ are Hilbert spaces, with $X_{1}$ and $Y_{1}$ continuously embedded in $X_{0}$ and $Y_{0}$, respectively, and $\mathcal{A}:X_{j}\to Y_{j}$ is a linear and topological isomorphism, for $j=0,1$, then, for $0<\theta<1$, $\mathcal{A}((X_{0},X_{1})_{\theta})=(Y_{0},Y_{1})_{\theta}$ and $\mathcal{A}:(X_{0},X_{1})_{\theta}\to(Y_{0},Y_{1})_{\theta}$ is a linear and topological isomorphism. This is immediate since $((X_{0},X_{1})_{\theta},(Y_{0},Y_{1})_{\theta})$ and $((Y_{0},Y_{1})_{\theta},(X_{0},X_{1})_{\theta})$ are, in the terminology of [13, §2], pairs of interpolation spaces relative to $(\overline{X},\overline{Y})$ and $(\overline{Y},\overline{X})$, respectively, where $\overline{X}=(X_{0},X_{1})$, $\overline{Y}=(Y_{0},Y_{1})$ (e.g., [3, Theorem 4.1.2]).
@@
We include the following corollary, although we will not use it subsequently, because it may be of independent interest. Note that the range of $s$ does not extend to s=(nd)/2s=-(n-d)/2 since, by [27, Thm 2.17], $H^{s}_{\Gamma}\neq\{0\}$ for s<(nd)/2s<-(n-d)/2, but $H^{-(n-d)/2}_{\Gamma}=\{0\}$ so that (e.g., [13, Theorem 2.2(iv)]) $\left(H_{\Gamma}^{s},H_{\Gamma}^{-(n-d)/2}\right)_{\theta}=\{0\}$ for all $s\in\mathbb{R}$ and $0<\theta<1$.
@@ 4 BVPs and BIEs
In this section we state the BVP and BIE that we wish to solve. We consider time-harmonic acoustic scattering of an incident wave $u^{i}$ propagating in $\mathbb{R}^{n+1}$ ($n=1,2$) by a planar screen $\Gamma$, a subset of the hyperplane $\Gamma_{\infty}=\mathbb{R}^{n}\times\{0\}$. We initially consider the case where $\Gamma$ is assumed simply to be non-empty and compact, for which a well-posed BVP/BIE formulation was presented in [16, §3.2]. We later specialise to the case where $\Gamma$ is a $d$-set for some $n-1<d\leq n$, and then further to the case where $\Gamma$ is a disjoint IFS attractor.
@@
The next result reformulates the BVP as a BIE. In this theorem $\mathcal{S}:H^{-1/2}_{\Gamma}\to C^{2}(D)\cap W^{1,{\rm loc}}(\mathbb{R}^{n+1})$ denotes the (acoustic) single-layer potential operator, defined by (e.g., [12, §2.2])
$\mathcal{S}\phi(x):=\langle\gamma^{\pm}(\sigma\Phi(x,\cdot)|_{U^{\pm}}),\overline{\phi}\rangle_{H^{1/2}(\Gamma_{\infty})\times H^{-1/2}(\Gamma_{\infty})},\quad x\in D,$ (32)
where $\overline{\phi}$ denotes the complex conjugate of $\phi$, [^15] If $\phi\in L_{2}(\Gamma_{\infty})\subset H^{-1/2}(\Gamma_{\infty})$, $\bar{\phi}$ is the usual complex conjugate, and this definition of the complex conjugate is extended to $H^{-1/2}(\Gamma_{\infty})$ by density. $\Phi(x,y):={\mathrm{e}}^{{\mathrm{i}}k|x-y|}/(4\pi|x-y|)$ ($n=2$), $\Phi(x,y):=({\mathrm{i}}/4)H^{(1)}_{0}(k|x-y|)$ ($n=1$), $H_{0}^{(1)}$ is the Hankel function of the first kind of order zero (e.g., [1, Equation (9.1.3)]), and $\sigma$ is any element of $C^{\infty}_{0,\Gamma}$ with $x\not\in\operatorname{supp}{\sigma}$. The $\pm$ in (32) indicates that either trace can be taken, with the same result. In the case when $\Gamma$ is the closure of a Lipschitz open subset of $\Gamma_{\infty}$ and $\phi\in L_{2}(\Gamma)$ the potential can be expressed as an integral with respect to (Lebesgue) surface measure, namely
$\mathcal{S}\phi(x)=\int_{\Gamma}\Phi(x,y)\phi(y)\,\mathrm{d}s(y),\qquad x\in D.$ (33)
@@
$S\phi:=P\gamma^{\pm}(\sigma\mathcal{S}\phi|_{U^{\pm}}),\qquad\phi\in H^{-1/2}_{\Gamma},$ (34)
@@
$a(\phi,\psi)=-\langle g,\psi\rangle_{H^{1/2}(\Gamma_{\infty})\times H^{-1/2}(\Gamma_{\infty})},\qquad\psi\in H^{-1/2}_{\Gamma}.$ (38)
@@
with $S^{\Omega}:\widetilde{H}{}^{-1/2}(\Omega)\to H^{1/2}(\Omega)$ the single-layer boundary integral operator on the Lipschitz screen $\Omega$, defined in the standard way (e.g., [11, §2.3]), so that
@@
for $\phi\in L_{2}(\Omega)$. The continuity and coercivity of $a^{\Omega}(\cdot,\cdot)$ (e.g., [11]) therefore implies the continuity and coercivity of $a(\cdot,\cdot)$, as the following lemma ([11], [12, §2.2]) states.
@@
Having computed $\phi$ by solving a Galerkin discretisation of (38), we will also evaluate $u(x)$ at points $x\in D$ using (35) and (32). Further, recall that (e.g., [10, Eqn. (2.23)], [32, p. 294])
@@
uniformly in $\hat{x}:=x/|x|$, where $u^{\infty}\in C^{\infty}(\mathbb{S}^{n})$ is the so-called far-field pattern of $u$. We will also compute this far-field pattern, given explicitly ([10, Eqn. (2.23)], [32, p. 294]) as
$u^{\infty}(\hat{x})=-\langle\gamma^{\pm}(\sigma\Phi^{\infty}(\hat{x},\cdot)),\overline{\phi}\rangle_{H^{1/2}(\Gamma_{\infty})\times H^{-1/2}(\Gamma_{\infty})},\quad\hat{x}\in\mathbb{S}^{n},$ (42)
@@
$\Phi^{\infty}(\hat{x},y):=\frac{{\mathrm{i}}k^{(n-2)/2}}{2(2\pi{\mathrm{i}})^{n/2}}\,\exp(-{\mathrm{i}}k\hat{x}\cdot y),\quad\hat{x}\in\mathbb{S}^{n},\;y\in\mathbb{R}^{n+1}.$ (43)
@@
The following lemma provides conditions under which a compact screen $\Gamma\subset\Gamma_{\infty}$ produces a non-zero scattered field. We note that a sufficient condition for $H^{-1/2}_{\Gamma}\neq\{0\}$ is that ${\rm dim_{H}}(\Gamma)>n-1$ [27, Thm 2.12], and that when $\Gamma$ is a $d$-set this is also a necessary condition [27, Thm 2.17].
Lemma 4.4 ([12, Thm 4.6]).

### v2

@@ 2.5 Function spaces on planar screens
For open sets $\Omega\subset\mathbb{R}^{n+1}$ (e.g. the exterior domain $D:=\mathbb{R}^{n+1}\setminus{\Gamma}$) we work with the classical Sobolev spaces [^8] Our notation follows [38]. Given an open set $\Omega\subset\mathbb{R}^{n+1}$, $W^{1}(\Omega)$ is the Sobolev space whose norm is defined “intrinsically”, via integrals over $\Omega$, while $H^{1}(\Omega)$ is defined (see §2.4) “extrinsically” as the set of restrictions to $\Omega$ of functions in $H^{1}(\mathbb{R}^{n+1})$. These spaces coincide if $\Omega$ is Lipschitz (e.g., [38, Theorem 3.16]), in particular if $\Omega=\mathbb{R}^{n+1}$, but not in general, in particular not, in general, for $\Omega=D$. $W^{1}(\Omega)$ and $W^{1}(\Omega,\Delta)$, normed by $\|u\|_{W^{1}(\Omega)}^{2}=\|u\|_{L_{2}(\Omega)}^{2}+\|\nabla u\|_{L_{2}(\Omega)}^{2}$ and $\|u\|_{W^{1}(\Omega,\Delta)}^{2}=\|u\|_{L_{2}(\Omega)}^{2}+\|\nabla u\|_{L_{2}(\Omega)}^{2}+\|\Delta u\|_{L_{2}(\Omega)}^{2}$ respectively, and their “local” versions $W^{1,{\rm loc}}(\Omega)$ and $W^{1,{\rm loc}}(\Omega,\Delta)$, defined as the sets of measurable functions on $\Omega$ whose restrictions to any bounded open $\Omega^{\prime}\subset\Omega$ are in $W^{1}(\Omega^{\prime})$ or $W^{1}(\Omega^{\prime},\Delta)$ respectively. We denote by $\gamma^{\pm}:W^{1}(U^{\pm})\to H^{1/2}(\Gamma_{\infty})$ and $\partial_{\mathrm{n}}^{\pm}:W^{1}(U^{\pm},\Delta)\to H^{-1/2}(\Gamma_{\infty})$ the standard Dirichlet and Neumann trace operators from the upper and lower half spaces $U^{\pm}:=\{x\in\mathbb{R}^{n+1},\pm x_{n+1}>0\}$ onto the hyperplane $\Gamma_{\infty}$, where the normal vector is assumed to point into $U^{+}$ in the case of the Neumann trace. Explicitly, the traces are the extension by density of $\gamma^{\pm}(u)(x):=\lim_{\begin{subarray}{c}x^{\prime}\to x\\ x^{\prime}\in U^{\pm}\end{subarray}}u(x^{\prime})$ and $\partial_{\mathrm{n}}^{\pm}u(x):=\lim_{\begin{subarray}{c}x^{\prime}\to x\\ x^{\prime}\in U^{\pm}\end{subarray}}\frac{\partial u}{\partial x_{n+1}}(x^{\prime})$ for $u\in C^{\infty}_{0}(\mathbb{R}^{n+1})|_{U^{\pm}}$ and $x\in\Gamma_{\infty}$. We note that if $u\in W^{1}(\mathbb{R}^{n+1})$ then $\gamma^{+}(u|_{U^{+}})=\gamma^{-}(u|_{U^{-}})$ [15, §2.1]. Finally, let $C^{\infty}_{0,\Gamma}$ denote the set of functions in $C^{\infty}_{0}(\mathbb{R}^{n+1})$ that equal one in a neighbourhood of $\Gamma$.
@@ 3 Wavelet decompositions
The spaces $\mathbb{H}^{t}(\Gamma)$ were defined in §2.4 for $\Gamma\subset\mathbb{R}^{n}$ a general $d$-set with $0<d\leq n$. In the case where $\Gamma$ is a disjoint IFS attractor (in the sense of §2.3, in which case, by Lemma 2.6, $d<n$), it was shown in [36] that, when $n-1<d={\rm dim_{H}}(\Gamma)<n$, the spaces $\mathbb{H}^{t}(\Gamma)$ can be characterised in terms of wavelet decompositions for t>0t>0.999More precisely, [36] showed that the Besov spaces $B^{p,q}_{\alpha}(\Gamma)$, as defined in [36, §6] (and see Definition A.1 below), can be characterised in this way for $\alpha>0$ and $1\leq p,q\leq\infty$, provided $\Gamma$ preserves Markov’s inequality, which holds in particular (see Remark A.6) if d>n1d>n-1. We show in Appendix A (see Corollary A.4 and Remark A.6) that, for $t>0$, $\mathbb{H}^{t}(\Gamma)=B^{2,2}_{t}(\Gamma)$ with equivalence of norms if $\Gamma$ is a $d$-set with d>n1d>n-1, so that this characterisation carries over to our trace spaces $\mathbb{H}^{t}(\Gamma)$. This, more precisely an extension of this characterisation to negative $t$, will be central to our BEM convergence analysis later. In the current section we recap the notation and main results from [36] that we will need, initially assuming only that $n-1<d\leq n$ and that the OSC is satisfied.
@@
These bases are hierarchical, in the sense that the basis for $\mathbb{W}_{\ell+1}$ contains that for $\mathbb{W}_{\ell}$. Furthermore, as noted in [36, p. 334] (and demonstrated in the proof of Theorem 5.1 below), elements of $\mathbb{L}_{2}(\Gamma)$ can be approximated arbitrarily well by elements of $\mathbb{W}_{\ell}$ as $\ell\to\infty$, which implies that
@@
The next result, which combines Theorems 1 and 2 in [36], provides a characterization of the space $\mathbb{H}^{t}(\Gamma)\subset\mathbb{L}_{2}(\Gamma)$ (introduced after (9)) for $n-1<d={\rm dim_{H}}(\Gamma)<n$ and $0<t<1$ in terms of the wavelet basis introduced above, under the assumption that $\Gamma$ is a disjoint IFS attractor. This assumption ensures that the piecewise-constant spaces $\mathbb{W}_{\ell}$ are contained in $\mathbb{H}^{t}(\Gamma)$ for all $t>0$ (see [36, Thm. 2]). In the statement of Theorem 3.1 the set $J_{\nu}$ is defined for $\nu\in\mathbb{Z}$ by
@@
Theorem 3.1 has the following important corollary, which is obtained by duality. In this corollary and subsequently (see, e.g., [16, Remark 3.8]), given an interval $\mathcal{I}\subset\mathbb{R}$ we will say that a collection of Hilbert spaces $\{H_{s}:s\in\mathcal{I}\}$, indexed by $\mathcal{I}$, is an interpolation scale if, for all $s,t\in\mathcal{I}$ and $0<\eta<1$, $(H_{s},H_{t})$ is a compatible couple (in the standard sense, e.g. [4, §2.3]) and if the interpolation space $(H_{s},H_{t})_{\eta}$ [^11] Here, and subsequently, $(H_{s},H_{t})_{\eta}$ denotes the standard complex interpolation space, $(H_{s},H_{t})_{[\eta]}$ in the notation of [4]; equivalently, the $K$- or $J$-method real interpolation spaces denoted $(H_{s},H_{t})_{\eta,2}$ in [4, 16], which are the same Hilbert spaces $(H_{s},H_{t})_{[\eta]}$, with equal norms, if the $K$- and $J$-methods are appropriately normalised (see [16, Remark 3.6] and [18]). coincides with $H_{\theta}$, for $\theta=(1-\eta)s+\eta t$, with equivalent norms. We will say that $\{H_{s}:s\in\mathcal{I}\}$ is an exact interpolation scale if, moreover, the norms of $(H_{s},H_{t})_{\eta}$ and $H_{\theta}$ coincide, for all $s,t\in\mathcal{I}$ and $0<\eta<1$.
@@
For (v) we note that by [16, Thm. 3.1] $\{\mathfrak{h}^{\tau}\}_{\tau\in\mathbb{R}}$ is an exact interpolation scale. The corresponding statement about $\{\mathbb{H}^{\tau}(\Gamma)\}_{-1<\tau<1}$, equipped with the norm $\|\cdot\|_{\tau}$, follows from Theorem 3.1 and parts (i)-(iii), combined with [^12] There is an inaccuracy in the statement of [16, Cor. 3.2]; the map $\mathcal{A}:\Sigma(\overline{H})\to\mathcal{Y}$ in that corollary needs to be injective as well as linear for the corollary to hold (see [18]). This injectivity follows automatically from the other conditions on $\mathcal{A}$ when (as in the application we make here) $H_{1}\subset H_{0}$, with continuous embedding, since then $\Sigma(\overline{H})=H_{0}$. [16, Cor. 3.2] and the fact that $\mathbb{H}^{0}(\Gamma)=\mathbb{L}_{2}(\Gamma)$ is unitarily isomorphic to $\mathfrak{h}^{0}$. Explicitly, in [16, Cor. 3.2], given $1<\tau_{0}<\tau_{1}<1$ we take $H_{j}=\mathbb{H}^{\tau_{j}}(\Gamma)$, $j=0,1$, $\mathcal{X}=\{0\}\cup\{(\nu,{\boldsymbol{m}},m):{\boldsymbol{m}}\in I_{\ell},\,\ell\in\mathbb{N}_{0},\,m\in\{1,\ldots,M-1\}\}$, $\mu$ to be the counting measure on $\mathcal{X}$, $\mathcal{A}$ to be the map taking $f\in\mathbb{H}^{\tau_{j}}(\Gamma)$ to the sequence $\boldsymbol{\beta}$ defined by (20) or (24) (as appropriate), and $w_{j}(0)=1$, $w_{j}((\nu,{\boldsymbol{m}},m))=2^{2\nu\tau_{j}}$. ∎
A basic interpolation result is that if $X_{0}\supset X_{1}$ and $Y_{0}\supset Y_{1}$ are Hilbert spaces, with $X_{1}$ and $Y_{1}$ continuously embedded in $X_{0}$ and $Y_{0}$, respectively, and $\mathcal{A}:X_{j}\to Y_{j}$ is a linear and topological isomorphism, for $j=0,1$, then, for $0<\theta<1$, $\mathcal{A}((X_{0},X_{1})_{\theta})=(Y_{0},Y_{1})_{\theta}$ and $\mathcal{A}:(X_{0},X_{1})_{\theta}\to(Y_{0},Y_{1})_{\theta}$ is a linear and topological isomorphism. This is immediate since $((X_{0},X_{1})_{\theta},(Y_{0},Y_{1})_{\theta})$ and $((Y_{0},Y_{1})_{\theta},(X_{0},X_{1})_{\theta})$ are, in the terminology of [16, §2], pairs of interpolation spaces relative to $(\overline{X},\overline{Y})$ and $(\overline{Y},\overline{X})$, respectively, where $\overline{X}=(X_{0},X_{1})$, $\overline{Y}=(Y_{0},Y_{1})$ (e.g., [4, Theorem 4.1.2]).
@@
We include the following corollary, although we will not use it subsequently, because it may be of independent interest. Note that the range of $s$ does not extend to s=(nd)/2s=-(n-d)/2 since, by [32, Thm 2.17], $H^{s}_{\Gamma}\neq\{0\}$ for s<(nd)/2s<-(n-d)/2, but $H^{-(n-d)/2}_{\Gamma}=\{0\}$ so that (e.g., [16, Theorem 2.2(iv)]) $\left(H_{\Gamma}^{s},H_{\Gamma}^{-(n-d)/2}\right)_{\theta}=\{0\}$ for all $s\in\mathbb{R}$ and $0<\theta<1$.
@@ 4 BVPs and BIEs
In this section we state the BVP and BIE that we wish to solve. We consider time-harmonic acoustic scattering of an incident wave $u^{i}$ propagating in $\mathbb{R}^{n+1}$ ($n=1,2$) by a planar screen $\Gamma$, a subset of the hyperplane $\Gamma_{\infty}=\mathbb{R}^{n}\times\{0\}$. We initially consider the case where $\Gamma$ is assumed simply to be non-empty and compact, for which a well-posed BVP/BIE formulation was presented in [19, §3.2]. We later specialise to the case where $\Gamma$ is a $d$-set for some $n-1<d\leq n$, and then further to the case where $\Gamma$ is a disjoint IFS attractor.
@@
The next result reformulates the BVP as a BIE. In this theorem $\mathcal{S}:H^{-1/2}_{\Gamma}\to C^{2}(D)\cap W^{1,{\rm loc}}(\mathbb{R}^{n+1})$ denotes the (acoustic) single-layer potential operator, defined by (e.g., [15, §2.2])
$\mathcal{S}\psi(x):=\langle\gamma^{\pm}(\sigma\Phi(x,\cdot)|_{U^{\pm}}),\overline{\psi}\rangle_{H^{1/2}(\Gamma_{\infty})\times H^{-1/2}(\Gamma_{\infty})},\quad x\in D,$ (30)
where $\overline{\psi}$ denotes the complex conjugate of $\psi$, [^14] If $\psi\in L_{2}(\Gamma_{\infty})\subset H^{-1/2}(\Gamma_{\infty})$, $\bar{\psi}$ is the usual complex conjugate, and this definition of the complex conjugate is extended to $H^{-1/2}(\Gamma_{\infty})$ by density. $\Phi(x,y):={\mathrm{e}}^{{\mathrm{i}}k|x-y|}/(4\pi|x-y|)$ ($n=2$), $\Phi(x,y):=({\mathrm{i}}/4)H^{(1)}_{0}(k|x-y|)$ ($n=1$), $H_{0}^{(1)}$ is the Hankel function of the first kind of order zero (e.g., [1, Equation (9.1.3)]), and $\sigma$ is any element of $C^{\infty}_{0,\Gamma}$ with $x\not\in\operatorname{supp}{\sigma}$. The $\pm$ in (30) indicates that either trace can be taken, with the same result. In the case when $\Gamma$ is the closure of a Lipschitz open subset of $\Gamma_{\infty}$ and $\psi\in L_{2}(\Gamma)$ the potential can be expressed as an integral with respect to (Lebesgue) surface measure, namely
$\mathcal{S}\psi(x)=\int_{\Gamma}\Phi(x,y)\psi(y)\,\mathrm{d}s(y),\qquad x\in D.$ (31)
@@
$S\psi:=P\gamma^{\pm}(\sigma\mathcal{S}\psi|_{U^{\pm}}),\qquad\psi\in H^{-1/2}_{\Gamma},$ (32)
@@
$a(\phi,\psi)=-\langle g,\psi\rangle_{H^{1/2}(\Gamma_{\infty})\times H^{-1/2}(\Gamma_{\infty})},\qquad\mbox{ for all }\psi\in H^{-1/2}_{\Gamma}.$ (36)
@@
with $S^{\Omega}:\widetilde{H}{}^{-1/2}(\Omega)\to H^{1/2}(\Omega)$ the single-layer boundary integral operator on the Lipschitz screen $\Omega$, defined in the standard way (e.g., [14, §2.3]), so that
@@
for $\phi\in L_{2}(\Omega)$. The continuity and coercivity of $a^{\Omega}(\cdot,\cdot)$ (e.g., [14]) therefore implies the continuity and coercivity of $a(\cdot,\cdot)$, as the following lemma ([14], [15, §2.2]) states.
@@
Having computed $\phi$ by solving a Galerkin discretisation of (36), we will also evaluate $u(x)$ at points $x\in D$ using (33) and (30). Further, recall that (e.g., [13, Eqn. (2.23)], [38, p. 294])
@@
uniformly in $\hat{x}:=x/|x|$, where $u^{\infty}\in C^{\infty}(\mathds{S}^{n})$ is the so-called far-field pattern of $u$ and $\mathds{S}^{n}$ is the unit sphere in $\mathbb{R}^{n+1}$. We will also compute this far-field pattern, given explicitly ([13, Eqn. (2.23)], [38, p. 294]) as
$u^{\infty}(\hat{x})=-\langle\gamma^{\pm}(\sigma\Phi^{\infty}(\hat{x},\cdot)),\overline{\phi}\rangle_{H^{1/2}(\Gamma_{\infty})\times H^{-1/2}(\Gamma_{\infty})},\quad\hat{x}\in\mathds{S}^{n},$ (40)
@@
$\Phi^{\infty}(\hat{x},y):=\frac{{\mathrm{i}}k^{(n-2)/2}}{2(2\pi{\mathrm{i}})^{n/2}}\,\exp(-{\mathrm{i}}k\hat{x}\cdot y),\quad\hat{x}\in\mathds{S}^{n},\;y\in\mathbb{R}^{n+1}.$ (41)
@@
The following lemma provides conditions under which a compact screen $\Gamma\subset\Gamma_{\infty}$ produces a non-zero scattered field. We note that a sufficient condition for $H^{-1/2}_{\Gamma}\neq\{0\}$ is that ${\rm dim_{H}}(\Gamma)>n-1$ [32, Thm 2.12], and that when $\Gamma$ is a $d$-set this is also a necessary condition [32, Thm 2.17].
Lemma 4.4 ([15, Thm 4.6]).